DOPED HELIUM CRYSTALS - THE PROMISES FOR MATRIX-ISOLATION SPECTROSCOPY

\underline{EUGENE B. GORDON}, Institute of Problems of Chemical Physics, Chernogolovka, Moscow Region, Russia 142432.

The novel universal technique for impurities embedding directly into solid He we recently elaborated ^{a}allows to acheive the guest particles concentrations as high as 10^{19} cm^{−3} (as small clusters suspended in a solid) and about 10^{16} cm^{−3} (as solitary isolated molecules or atoms) under a sample growth rate up to 3 mm per minute. Contrary to the liquid He droplets technique we can investigate the temperature and pressure dependecies, whereas the procedure of a sample preparation and replacement is more simple than for matrix-isolation by p-H_{2}. The promises for spectroscopic applications of the new matrix and the first experimental results will be discussed .

aE. B. Gordon, G. Frossati, A. Usenko. Phys. Rev. Lett.(submitted)